# Parrot: Pareto-optimal Multi-Reward Reinforcement Learning Framework for Text-to-Image Generation

### v1

@@ 5.1 Experiment Setting
Figure 3: Comparison of Parrot and diffusion-based RL baselines. From left to right, we provide results of Stable diffusion 1.5 [38] (1st column), DPOK [10] (2nd column) with the weighted sum, the Promptist [13] (3rd column), and the Parrot (4th column).
@@ 5.2 Qualitative Analysis
Comparison with Baselines: Fig 3 shows the visual comparison of Parrot and multiple baselines. We include results from Stable Diffusion 1.5, DPOK [10] with a weighted sum of rewards, Promptist [13], and Parrot. DPOK exclusively fine-tunes the T2I model, while Promptist focuses on fine-tuning only the prompt expansion network. Parrot shows visually better images, particularly in aspects like color combination, cropping, perspective, and fine details in the image. This improvement can be attributed to Parrot’s T2I model being fine-tuned together with the prompt expansion model that incorporates aesthetic keywords during training. Parrot generates results that are more closely aligned with the input prompt, as well as more visually pleasing.
Fig. 4 shows the training curve comparison of Parrot and using a linear combination of the reward scores. Each subgraph represents a reward. WS1 and WS2 denote two different weights with multiple reward scores. WS1 places greater emphasis on the aesthetic score, while WS2 adopts balanced weights across aesthetics, human preference, text-image alignment, and image sentiment. Employing the weighted sum of multiple rewards leads to a decrease in the image sentiment score, despite notable enhancements in aesthetics and human preference. In contrast, Parrot consistently exhibits improvement across all metrics.
@@
TIA ($\uparrow$)
Aesth ($\uparrow$)
HP ($\uparrow$)
IS ($\uparrow$)
SD 1.5 [38]
0.2322
0.5755
0.1930
0.3010
DPOK [10] (WS)
0.2337
0.5813
0.1932
0.3013
Parrot w/o PE 0.2355
0.6034
0.2009
0.3018
Promptist [13]
0.1449
0.6783
0.2759
0.2518
Parrot
0.1667
0.7396 0.3411 0.3132
Table 1: Quantitative comparison between Parrot and alternatives on the Parti dataset [53]. Abbreviations: WS - Weighted Sum; PE - Prompt Expansion; TIA - Text-Image Alignment; Aesth - Aesthetics; HP - Human Preference; IS - Image Sentiment. TIA score is measured against the original prompt without expansion.
@@ 5.3 Quantitative Evaluation
Comparison with Baselines: Table 1 presents our results of the quality score across four quality rewards: text-image alignment score, aesthetic score, human preference score, and emotion score. The first group shows methods without prompt expansion, and the second group compares methods with expansion. The prompt expansion and T2I generation are performed on the PartiPrompts [53]. Using a set of 1631 prompts, we generate 32 images for each text input and calculate the average for each metric. “w/o Prompt Expansion” indicates the generation of images solely based on original prompt without expansion. We compare with DPOK [10] with a weighted sum (WS) of rewards, using balanced weights of $\{0.25,0.25,0.25,0.25\}$. For Promptist [13], we generate prompt expansion from their model. Our method outperforms both compared methods in aesthetics, human preference and sentiment scores. The text-image alignment score is measured with the original prompt before expansion for fair comparison. As a result, the group without prompt expansion generally shows a higher text-image alignment score. Parrot shows better text-image alignment in each subgroup.
@@
Aesthetics “Which image shows better aesthetics without blurry texture, unnatural focusing, and poor color combination?”
Human Preference “Which generated image do you prefer?”
Text-Image Alignment “Which image is well aligned with the text?”
Image Sentiment “Which image is closer to amusement, excitement, and contentment?”
@@
User Study: We conduct a user study using MTurk [1] with generated images from 100 random prompts in the PartiPrompts [53]. Five models are compared: Stable Diffusion v1.5, DPOK [10] with an equal weighted sum, Promptist [13], Parrot without prompt expansion, and Parrot. Each rater is presented with the original prompt (before expansion) and a set of five generated images, with the image order being randomized. Raters are then tasked with selecting the best image from the group, guided by questions outlined in Table 2. Each question pertains to a specific quality aspect: aesthetics, human preference, text-image alignment, and image sentiment. For each prompt, 20 rounds of random sampling are conducted and sent to different raters. The aggregated user study results, illustrated in Fig 5, show that Parrot significantly outperforms other baselines across all dimensions.
@@
Figure 7: The comparisons of the diffusion fine-tuning between Pareto-optimal multi-reward RL and single reward RL. We show results with same seed from various methods: Stable Diffusion 1.5 [38] (1st column), T2I model fine-tuned with the aesthetic model (2nd column), the human preference model (3rd column), text-image alignment (4th column), image sentiment (5th column), Parrot without Pareto-optimal selection (6th column) and Parrot (7th column). Parrot is effective to generate acceptable images without sacrificing one of quality signals. For example, T2I model fine-tuned with a single quality signal such as aesthetics, human preference and image sentiment results in text-image misalignment. In contrast, our method achieves a balanced visual outcome across multiple criteria.
Figure 8: Results of original prompt centered guidance. As we expand the prompt, the content in the generated image often fades away (2nd column). We observe that this guidance is helpful for keeping the main content of the original prompt.
@@ 5.4 Ablations
Effect of Pareto-optimal Multi-reward RL: To show the efficacy of Pareto-optimal Multi-reward RL, we conduct an ablation study by removing one reward model at a time. Fig 6 shows quantitative results using one hundred random text prompts from the Promptist [13]. We observe that our training scheme improves multiple target objectives. Fig 7 shows the visual comparison between Parrot, Parrot with a single reward, and Parrot without selecting the batch-wise Pareto-optimal solution. Using a single reward model tends to result in degradation of another reward, especially text-image alignment. For example, in the third column, results of the first row miss the prompt a tophat, even though the Stable Diffusion result includes that attribute. On the other hand, Parrot results capture all prompts, improving other quality signals, such as aesthetics, image sentiment and human preference.
Effect of Original Prompt Centered Guidance: Fig 8 shows the effect of the proposed original prompt-centered guidance. As evident from the figure, utilizing only the expanded prompt as input often results in the main content being overwhelmed by the added context. For instance, given the original prompt “A shiba inu”, the result from the expanded prompt shows a zoomed-out image and the intended main subject (shiba inu) becomes small. The proposed original prompt-centered guidance effectively addresses this issue, generating an image that faithfully captures the original prompt while incorporating visually more pleasing details.
@@ 6 Conclusion and Limitation
In this paper, we propose Parrot, a framework designed to improve T2I generation by effectively optimizing multiple quality rewards using RL. With batch-wise Pareto-optimal selection, we effectively balance the optimization process of multiple quality rewards, improving each quality metric. Furthermore, through co-training the T2I model and the prompt expansion model, Parrot can generate higher-quality images. Additionally, our original prompt centered guidance technique enables the control over the persistence of the main content in user input, ensuring that the generated image remains faithful to the user prompt. Results from the user study indicate that Parrot significantly elevates the quality of generated images across multiple criteria including text-image alignment, human preference, aesthetics, and image sentiment. While Parrot has demonstrated effectiveness in enhancing generated image quality, its efficacy is limited by the quality metrics it relies on. Consequently, advancements of the generated image quality metrics will directly enhance the capabilities of Parrot. Additionally, Parrot is adaptable to a broader range of rewards that quantify generated image quality.
@@
  • Sec. B: more ablation studies on better tradeoff for different rewards.
  • Sec. C: more ablation studies on original prompt guidance and training scheme of Parrot.
  • Sec. D: more visual examples to show the advancements of Parrot.
@@ A. Implementation Details
B. Better Tradeoff for Different Rewards
To verify that Parrot achieved better tradeoff for different rewards, we generate 1000 images from common animal dataset [4], where each text prompt consists of the name of a common animal. As shown in Fig. 9, using only text-image alignment with reward-specific preference $<$Reward 3$>$ yields images with higher text-image alignment score, while using only aesthetic model with reward-specific preference $<$Reward 1$>$ yields images with higher aesthetic score. In the case of using two reward-specific preferences $<$Reward 1$>$, $<$Reward 3$>$, we observe that scores are balanced and show that better results across multiple rewards compared with Parrot without Pareto optimal selection.
Figure 9: Comparing Parrot with/without Pareto Optimal. When we apply Pareto-optimal selection, we observe that each reward score increases based on reward-specific preference.
C. Prompt Guidance and Training Scheme
@@
D. More Visual Examples

### v2

@@ 5.1 Experiment Setting
Figure 3: Comparison of Parrot and diffusion-based RL baselines. From left to right, we provide results of Stable diffusion 1.5 [39] (1st column), DPOK [12] (2nd column) with the weighted sum, Promptist [15] (3rd column), and Parrot (4th column).
@@ 5.2 Qualitative Analysis
Comparison with Baselines: Fig. 3 shows the visual comparison of Parrot and multiple baselines. We include results from Stable Diffusion 1.5, DPOK [12] with a weighted sum of rewards, Promptist [15], and Parrot. DPOK exclusively fine-tunes the T2I model, while Promptist focuses on fine-tuning only the prompt expansion network. Parrot shows visually better images, particularly in aspects like color combination, cropping, perspective, and fine details in the image. This improvement can be attributed to Parrot’s T2I model being fine-tuned together with the prompt expansion model that incorporates aesthetic keywords during training. Parrot generates results that are more closely aligned with the input prompt, as well as more visually pleasing.
Weighted sum vs. Parrot: Fig. 4 shows the training curve comparison of Parrot and using a linear combination of the reward scores. Each subgraph represents a reward. WS1 and WS2 denote two different weights with multiple reward scores. WS1 places greater emphasis on the aesthetic score, while WS2 adopts balanced weights across aesthetics, human preference, text-image alignment, and image sentiment. Employing the weighted sum of multiple rewards leads to a decrease in the image sentiment score, despite notable enhancements in aesthetics and human preference. In contrast, Parrot consistently exhibits improvement across all metrics.
@@
TIA ($\uparrow$) Aesth. ($\uparrow$) HP ($\uparrow$) Sent. ($\uparrow$) Average ($\uparrow$)
SD 1.5 [39] 0.2322 0.5755 0.1930 0.3010 0.3254
DPOK [12] (WS) 0.2337 0.5813 0.1932 0.3013 0.3273 (+0.58%)
Parrot w/o PE 0.2355 0.6034 0.2009 0.3018 0.3354 (+3.07%)
Parrot T2I Model Tuning Only 0.2509 0.7073 0.3337 0.3052 0.3992 (+22.6%)
Promptist [15] 0.1449 0.6783 0.2759 0.2518 0.3377 (+3.77 %)
Parrot with HP Only 0.1543 0.5961 0.3528 0.2562 0.3398 (+4.42 %)
Parrot PEN Tuning Only 0.1659 0.6492 0.2617 0.3131 0.3474 (+6.76 %)
Parrot w/o Joint Optimization 0.1661 0.6308 0.2566 0.3084 0.3404 (+4.60 %)
Parrot w/o ori prompt guidance 0.1623 0.7156 0.3425 0.3130 0.3833 (+17.8 %)
Parrot 0.1667 0.7396 0.3411 0.3132 0.3901 (+19.8 %)
Table 1: Quantitative comparison between Parrot and alternatives on the Parti dataset [53]. Abbreviations: WS - Weighted Sum; PE - Prompt Expansion; TIA - Text-Image Alignment; Aesth. - Aesthetics; HP - Human Preference; Sent. - Image Sentiment. TIA score is measured against the original prompt without expansion.
@@ 5.3 Quantitative Evaluation
Comparison with Baselines: Table 1 presents our results of the quality score across four quality rewards: text-image alignment score, aesthetic score, human preference score, and emotion score. The first group shows methods without prompt expansion, and the second group compares methods with expansion. The prompt expansion and T2I generation are performed on the PartiPrompts [53]. Using a set of 1632 prompts, we generate 32 images for each text input and calculate the average for each metric.
w/o PE indicates the generation of images solely based on original prompt without expansion. Note that w/o PE is not the same as Parrot T2I Model Tuning Only. The former takes a model trained with PEN and removes it during inference, which results in a notable disparity between prompts used in training and testing. The latter trains with the multi-objective RL using only the T2I model, and it shows substantial improvement upon DPOK [12] (weighted sum) with balanced weights of $\{0.25,0.25,0.25,0.25\}$. This shows that the proposed RL method is indeed effective in multi-objective optimization.
For Promptist [15], we generate prompt expansion from their model. Parrot with HP Only shows solely using a human prefernce reward leads to decline in others. Parrot PEN Tuning Only shows suboptimal aesthetics and human preference. Parrot w/o Joint Optimization shows suboptimal results than Parrot which demonstrates the necessity of jointly optimizing PEN and T2I models.
Our method outperforms both compared methods in aesthetics, human preference and sentiment scores. The text-image alignment score is measured with the original prompt before expansion for fair comparison. As a result, the group without prompt expansion generally shows a higher text-image alignment score. Parrot shows better text-image alignment in each subgroup.
@@
Aesthetics ‘‘Which image shows better aesthetics without blurry texture, unnatural focusing, and poor color combination?’’
Human Preference ‘‘Which generated image do you prefer?’’
Text-Image Alignment ‘‘Which image is well aligned with the text?’’
Image Sentiment ‘‘Which image is closer to amusement, excitement, and contentment?’’
@@
User Study: We conduct a user study using MTurk [1] with generated images from 100 random prompts in the PartiPrompts [53]. Five models are compared: Stable Diffusion v1.5, DPOK [12] with an equal weighted sum, Promptist [15], Parrot without prompt expansion, and Parrot. Each rater is presented with the original prompt (before expansion) and a set of five generated images, with the image order being randomized. Raters are then tasked with selecting the best image from the group, guided by questions outlined in Table 2. Each question pertains to a specific quality aspect: aesthetics, human preference, text-image alignment, and image sentiment. For each prompt, 20 rounds of random sampling are conducted and sent to different raters. The user study results, illustrated in Fig. 6, show that Parrot outperforms other baselines across all dimensions.
Proportion of non-dominated points: Using batch size of 256, we observe that in a batch around $20\%$ to $30\%$ are non-dominated points and the proportion of non-dominated points in single batch slightly increase as training proceeds.
Figure 7: Comparing Parrot with/without Pareto Optimal. By prepending different reward-specific preferences, Parrot can change the tradeoff between rewards. Specifically, “<reward 1>”, “<reward 2>”, “<reward 3>”, “<reward 4>” indicates aesthetics, human preference, text-image alignment, and image sentiment respectively. With Pareto-optimal selection, each reward increases based on reward-specific preference.
@@ 5.4 Ablations
Figure 8: The comparisons of the diffusion fine-tuning between Pareto-optimal multi-reward RL and single reward RL. We show results with same seed from various methods: Stable Diffusion 1.5 [39] (1st column), T2I model fine-tuned with the aesthetic model (2nd column), the human preference model (3rd column), text-image alignment (4th column), image sentiment (5th column), Parrot without Pareto-optimal selection (6th column) and Parrot (7th column). Parrot is effective to generate acceptable images without sacrificing one of quality signals. For example, T2I model fine-tuned with a single quality signal such as aesthetics, human preference and image sentiment results in text-image misalignment, while our method achieves a balanced visual outcome across multiple criteria.
Effect of Pareto-optimal Multi-reward RL: To show the efficacy of Pareto-optimal Multi-reward RL, we conduct an ablation study by removing one reward model at a time. Fig. 6 shows quantitative results using one hundred random text prompts from the Promptist [15]. We observe that our training scheme improves multiple target objectives.
To verify whether Parrot achieved better trade-off for different reward scores, we generate 1000 images from common animal dataset [4], where each text prompt consists of the name of a common animal. As shown in Fig. 7, using only text-image alignment with reward-specific preference $<$reward 3$>$ generates images with higher text-image alignment score, while using only aesthetic model with reward-specific preference $<$reward 1$>$ yields images with higher aesthetic score. In the case of using two reward-specific preferences $<$reward 1$>$, $<$reward 3$>$, we observe that scores are balanced and show that better results across multiple rewards than Parrot without Pareto optimal selection.
Fig 8 shows the visual comparison between Parrot, Parrot with a single reward, and Parrot without selecting the batch-wise Pareto-optimal solution. Using a single reward model tends to result in degradation of other rewards, especially text-image alignment. In the third column, results of the first row miss the text a tophat in input prompt, even though the Stable Diffusion result includes that attribute. On the other hand, Parrot results capture all prompts, improving other quality signals, such as aesthetics, image sentiment and human preference.
Effect of Original Prompt Centered Guidance: Fig 9 shows the effect of the proposed original prompt-centered guidance. As evident from the figure, using only the expanded prompt as input often results in the main content being overwhelmed by the added context. For instance, given the original prompt “A shiba inu", the result from the expanded prompt shows a zoomed-out image and the intended main subject (shiba inu) becomes small. The proposed original prompt-centered guidance effectively addresses this issue, generating an image that faithfully captures the input prompt while incorporating visually more pleasing details.
Figure 9: Results of original prompt centered guidance. As we expand the prompt, the content in the generated image often fades away. This guidance is helpful for keeping the main content of the original prompt.
@@ 6 Conclusion and Limitation
We propose Parrot, a novel multi-reward optimization algorithm aimed to improve text-to-image generation by effectively optimizing multiple quality rewards using RL. With batch-wise Pareto-optimal selection, Parrot adaptively balance the optimization of multiple quality rewards. By applying Parrot to jointly finetune both the T2I model and the prompt expansion model, we achieve the generation of higher-quality images with richer details. Additionally, our original prompt centered guidance technique ensures that the generated image maintains fidelity to the user prompt after prompt expansion during inference. Results from the user study indicate that Parrot significantly improves the quality of generated images across multiple criteria, including text-image alignment, human preference, aesthetics, and image sentiment. While Parrot has shown effectiveness in enhancing generated image quality, its efficacy is limited by the quality metrics it relies on. Therefore, advancements of the generated image quality metrics will directly enhance the capabilities of Parrot. Additionally, Parrot is adaptable to a broader range of rewards that quantify generated image quality.
@@
Supplementary Material
Parrot: Pareto-optimal Multi-Reward
Reinforcement Learning Framework for
Text-to-Image Generation
@@
  • Sec. B: more ablation studies on original prompt guidance and training scheme of Parrot.
  • Sec. C: more visual examples to show the advancements of Parrot.
@@ A. Implementation Details
B. Prompt Guidance and Training Scheme
@@
C. More Visual Examples